A particle moves in a circle of radius 20 cm with a linear speed of 10 m/s . Find the angular velocity.

Grade:12th pass

2 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The angular velocity is given by:
\omega =\frac{v}{r}=\frac{10}{20}=0.5 rad/sec
Regards
Sumit
Apoorva Arora IIT Roorkee
9 years ago
As we know that v=wr,where
  • v is the linear velocity
  • w is the angular velocity
  • r is the radius of the circular path
So the angular velocity is
\omega = v/r
\omega = 10/20= 50 rad/s.